OpenStudy (anonymous):

The height of an ostrich is 33 inches more than 3 times the height of a kiwi, when the height of an ostrich is 162 inches. Write and solve an equation to find the height of a kiwi. Show your answer is reasonable.

OpenStudy (anonymous):

okay, so let's say that the height of an ostrich is represented by the letter O, and the height of the kiwi is represented by the letter K. from what the question says, the height of an ostrich is 33 inches more than 3 times the height of the kiwi. that means, (height of ostrich) = 33 + 3*(height of kiwi) \[O = 33+3K\] we know that the ostrich is 162 inches tall, so we plug that into the equation. \[162 = 33+3K\] we solve to find K in this series of equations: \[162-33 = 3K\] \[129 = 3K\] \[129/3 = K\] \[43 = K\] there we go! that is the height of a kiwi. XD